A firm in competitive market receives $600 in total revenue and has $10 as marginal revenue. 

What will be the average revenue and how many units were sold?


1
Expert's answer
2021-06-29T08:45:28-0400


Total Revenue TR =$600

Marginal Revenue MR = $10


"\frac{d(TR)}{dQ}=MR ==>\frac{d(PQ)}{dQ}=MR"


as "P=MR"

"P= {}"$"10"

In perfect competition MR = Average Revenue.

As average revenue =$10

And "Q=\frac{TR}{P}=\frac{600}{10}=60"